OpenStudy (anonymous):

let \[F(r) = \int\limits_{0}^{x}e ^{-t^2}dt\] a) find Maclorian Series (i.e Taylor Series about 0) of F(r). (You may use Maclorian Series of well-known functions) b) approximate F(1/2) with error less then 0.00001

OpenStudy (anonymous):

ahh finally someone :D

OpenStudy (anonymous):

i got something but i have go afk for like 20-30 mins ill say what i found :D when i back

OpenStudy (anonymous):

So it goes 1-t^2 +t^4/2 -t^6/6 etc 1,3/4,13/16,51/64 so maybe u can use Taylor's Theorem/Remainder to calculate error bounds until u are within your limit...

OpenStudy (anonymous):

solve my question mybe ...

OpenStudy (anonymous):

I don't do numbers..:-)

OpenStudy (anonymous):

well :D let me write what i got

OpenStudy (anonymous):

\[e^t = \sum_{n=0}^{\infty}(t^n / n!) \nu t \epsilon iR\]

OpenStudy (anonymous):

What's that?

OpenStudy (anonymous):

and hence; \[e ^{-t^2} = \sum_{n=0}^{\infty}((-ı^n)t ^{2n})/n!\]

OpenStudy (anonymous):

\[\int\limits_{0}^{x}e^{-t^2}dt = \sum_{n=0}^{infity}((-1)^n x^{2n+1}) / ((2n+2)n!) = F(x)\]

OpenStudy (anonymous):

I already did that above (and F(1/2))

OpenStudy (anonymous):

\[\nu x \epsilon ı\mathbb{R}\]

OpenStudy (anonymous):

What's that say?

OpenStudy (anonymous):

thats what i got :D

OpenStudy (anonymous):

dunno :P just calculated what i see there o.O

OpenStudy (anonymous):

So it goes 1-t^2 +t^4/2 -t^6/6 etc Is the expansion 1,3/4,13/16,51/64 Is F(1/2) Now u have to test error.

OpenStudy (anonymous):

not calculate we can say order

OpenStudy (anonymous):

yeap, but i dunno how to test error -.-'

OpenStudy (anonymous):

It's Taylor's Theorem..if remainder = f(x) -p(x) then remainder = f^(n+1)*c/(n+1)! * x^(n+1) and u have to calculate function and each partial to test for your limit...

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Taylor%27s_Theorem

OpenStudy (anonymous):

o.O hmm

OpenStudy (anonymous):

Don't look at me:-)

OpenStudy (anonymous):

lol :) really good stuff

OpenStudy (anonymous):

You are calculating an "error" against a function which is only an approximation to begin with, lol.

OpenStudy (anonymous):

nono i mean the theorem ^.^

OpenStudy (anonymous):

\[f(1/2) \approx (1/2) - (1/24) + (1/320) - (1/5376)\]

OpenStudy (anonymous):

?? is this ture? :D

OpenStudy (anonymous):

Here http://www.wolframalpha.com/input/?i=series+e^-x^2

OpenStudy (anonymous):

hmm